OpenStudy (anonymous):

Given: -Square ABCD with side length 2. -Two quarter circles, one with center A and the other at center B. Call their intersection E. What is the area enclosed by Arc CE, Arc DE, and side CD? I've already done the problem, but I want to verify my answer. Also, my solution does not use calculus, trigonometry, and any other form of high-powered math. The whole solution only required geometry.

OpenStudy (amistre64):

ack!!...picture would help

OpenStudy (anonymous):

No problem. Give me a minute.

OpenStudy (anonymous):

This ought to help. Essentially, finding the area of the shaded.

OpenStudy (amistre64):

thats easier :)

OpenStudy (amistre64):

we essentially got 4 areas, lets call them a,b,c, and d

OpenStudy (anonymous):

When I did that, I formuated the equation a=4-b-c-d

OpenStudy (amistre64):

thats good... we also need to realize that: b+d = c+d

OpenStudy (amistre64):

b = c

OpenStudy (amistre64):

a = 4-2b-d right?

OpenStudy (amistre64):

lets try to fill in some values here: lets say b = 5, c = 5, and d = 10 we know that: (5+10) + (5+10) but the area here is exagerated by "an extra 10"

OpenStudy (amistre64):

we have: the true area for these overlapping circles is: 2b + 2d - d right?

OpenStudy (amistre64):

2b + d jsut gets us right back to where we were .... can we use the circumference of a circle in our solution?

OpenStudy (amistre64):

2pi(2) = 4pi for a whole circle; which means that pi is the length for a quarter of it......I think I got it. unless you already know the answer :)

OpenStudy (amistre64):

do we know that area of an equilateral triangle?

OpenStudy (amistre64):

that middle triangle has an area of: sqrt(3)

OpenStudy (anonymous):

Thats what I did too!

OpenStudy (amistre64):

each side section has an area of 1 so the total unshaded area is 2sqrt(3)

OpenStudy (amistre64):

pi(r^2) is the area of a complete circle the r here =2 4pi is the total area of our representative circle... we only want 30 out of 360 of that circle

OpenStudy (amistre64):

30 degrees is 1/6 of a circle right?

OpenStudy (amistre64):

so the area of each side should be 4pi/6 = 2pi/3

OpenStudy (amistre64):

got something turned around...hold on.

OpenStudy (amistre64):

360/30 = 12, so we have 1/12 of a circles area.... 4pi(1/12) = 4pi/12 = pi/3... am i right?

OpenStudy (amistre64):

it should be 2pi/3 for the other stuff

OpenStudy (amistre64):

a =4-sqrt(3)-(2pi/3) a = 12 - 3sqrt(3) - 2pi ---------------- 3 is my answer

OpenStudy (amistre64):

a = .1735 if I did it right :)

OpenStudy (anonymous):

I didn't evaluate for that, but the formula for a is what I got too! Awesome, thanks! You just earned a fan.